OpenStudy (anonymous):

Combine like terms. 8 + 16r – 9r A. 8 + 7r B. 15r C. 15 D. 8 – 7r

OpenStudy (anonymous):

A

OpenStudy (a1234):

16r - 9r = 7r So, 8 + 7r.

OpenStudy (anonymous):

Add and subtract just like you would normally, with same variables.

OpenStudy (anonymous):

16r-9r=7r

OpenStudy (anonymous):

A just think as in the numbers with "r" are together. and basically a regular equation but just with a variable

OpenStudy (anonymous):

the answer is A.

OpenStudy (anonymous):

thanks all of you i wish i could give you all metals!

OpenStudy (luigi0210):

If it helps. get rid of the variable and replace it with x.. \[\LARGE 8+(16x-9x)\] Remeber tho \(x=r\)